Steam engine 22 and engine 65 are stationed at different locations along the Oldbridge line 86 miles apart. Engine 22 leaves the station at 1:00. Due to a communication error, engine 65 leaves the station at 2:00, heading toward engine 22. If engine 22 travels at a constant rate of 30 mph and Engine 65 travels at a constant rate of 40mph, by what time will the signalman have to switch one of the engines to a side track before they collide.
Answer by checkley77(12844)

D=RT
D=30T FOR E22
D=40(T-1) FOR E65
30T+40(T-1)=86
30T+40T-40=86
70T=86+40
70T=126/70
T=126/70
T=1.8 HOURS THEY WOULD MEET.
1:00+1:48=2:48 IS THE TIME OF THE SWITCH.